< Prev Next Question 12 sineering Corporation purchased comveyor equipment with a ist price of $9.300 Presented below are three independent cases related to the equipmgnt ) Avora aid ash for the equipment B days ater h h ndsre e10,n/30.Assume that caulomant purchases re iil (b) Ayayai traded in equipment with a book value of $ 2.200 (initial cost $ 7,300),and paid $ 9,700 in cash one month after the purchase. The old equipment could have gross been sold for $400 at the date of trade.(The exchange has commercial substance.) (el) Ayayai gave the vendor a $ 11,900 zero-interest-bearing note for the equipment on the date of purchase. The note was duein one year and was paid on time As that the effective-interest rate in the market was 9% Prepare the general journal entries required to record the acquisition and payment in each of the independent cases above.(Round decimal places, eg 1.25124 and final answers to O decimal places, e s 5,275.Credit account tit entry is required, select No Entry for the account titles and enter O for the amounts) present value factor calculations to 5 les are automatically indented when amount is entered. Do not indent manually fn No. Account Titles and Explanation Debit Credit
Answer #1

a) Equipment Dr. 9300

Accounts Payable Cr  9300

(To record the purchase of equipment on account)

Accounts Payable Dr.  9300

Cash Cr 9114

Equipment Cr. 186

(To record the payment on account)

b) Equipment(new) Dr. 10100

Loss on disposal of equipment Dr 1800

Accumulated Depreciation-equipment Dr. 5100

Accounts Payable Cr. 9700

Equipment (Old) Cr. 7300

(To record the purchase of equipment on account)

Accounts Payable Dr.  9700

Cash Cr 9700

(To record the payment on account)

c) Equipment Dr. 10829

Discount on Notes Payable Cr. 1071

Notes Payable Cr. 11900

(To record the purchase of equipment with a note)

Interest Expense Dr. 1071

Notes Payable Dr. 11900

Discount on Notes Payable Cr. 1071

Cash Cr. 11900

(To Record the payment of the note)

Notes:

a) Equipment=$9300 * 2%=183

b) Accumulated Depreciation = $7300-2200=$5100

Cost=7300

Less:Accumulated Depreciation=5100

Book Value of Equipment (Old)=2200

Less: Fair Value of Equipment (Old)=400

Loss on disposal of equipment=1800

Cost(9700+400)=10100

c) Equipment = 11900*91%=10829

  Discount on Notes Payable= 11900-10829=1071